OpenStudy (anonymous):

A positive integer and a negative integer are 18 units apart. The absolute value of the positive integer is 5 times the absolute value of the negative integer

OpenStudy (anonymous):

\[\left| a\right|=5\left| b \right|\]  So to find the answer we are looking for \[\left| a \right|+(5\left| b \right|)=18\]